OpenStudy (anonymous):

Find a unit vector with positive first coordinate that is orthogonal to the plane through the points P = (-2, 2, -3), Q = (3, 7, 2), and R = (3, 7, 5).

OpenStudy (anonymous):

Use the points to find two vectors, take the cross product of those vectors to get an orthogonal vector to the plane. Then normalize that vector.

OpenStudy (anonymous):



OpenStudy (anonymous):

So first find \(v\) and \(w\). Then find \(u=v\times w\). Then normalize \(u\) with \(u/\|u\|\) Where \(\|u\| = \sqrt{u\cdot u}\)

OpenStudy (anonymous):

@theone5526 What is left that you don't get?

OpenStudy (anonymous):

@wio one second.. working through what youve given me right now

OpenStudy (anonymous):

I got it thanks!

OpenStudy (anonymous):

Prodigious.